1. (x2-4xy-2y2)dx + (y2-4xy-2x2)dy=0
  2. (x2-xtan2y+sec2y)dy = (tany-2xy-y)dx
  3. (2xy4ey+2xy3+y)dx + (x2y4ey-x2y2-3x)dy=0
  4. (y/x secy-tany)dx+(secy logx-x)dy=0
  5. y dy/dx +4x/3 -y2/3x =0
1
Expert's answer
2022-01-13T07:54:37-0500

1.

"M=x^2-4xy-2y^2"

"N=y^2-4xy-2x^2"

"\partial M/\partial y=-4x-4y=\partial N/\partial x"

so, equation is exact


"\int Mdx=\int (x^2-4xy-2y^2)dx=x^3/3-2x^2y-2y^2x"


"\int Ndy=\int (y^2-4xy-2x^2)dx=y^3/3-2x^2y-2y^2x"


omitting -2xy2 - 2x2y which already occur in ∫Mdx:

"x^3\/3-2x^2y-2y^2x+y^3\/3=c"


2.

"M= -tany+2xy+y"

"N=x^2-xtan^2y+sec^2y"

"\partial M/\partial y=2x-tan^2y=\partial N/\partial x"

so, equation is exact


"F=\int Mdx=\int ( -tany+2xy+y)dx=x^2y+xy-xtany+g(y)"


"\partial F/\partial y=x^2+x-xsec^2y+g'(y)"

"x^2+x-xsec^2y+g'(y)=N=x^2-xtan^2y+sec^2y"

"g'(y)=sec^2y"

"g(y)=\int sec^2y dy=tany +c"

"F=x^2y+xy-xtany+tany+c=0"


3.

"M=2xy^4e^y+2xy^3+y"

"N=x^2y^4e^y-x^2y^2-3x"

"\frac{\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x}}{M}=\frac{-8xy^2-4-8xy^3e^y}{2xy^4e^y+2xy^3+y}=-\frac{4}{y}"


Integrating factor:

"e^{\int(-4/y)dy}=1/y^4"


then:

"\frac{1}{y^4}(2xy^4e^y+2xy^3+y)dx +\frac{1}{y^4} (x^2y^4e^y-x^2y^2-3x)dy=0"


"M_1=2xe^y+2x\/y+1\/y^3"

"N_1=x^2e^y-x^2\/y^2-3x\/y^4"


"\int M_1dx+\intop"(terms of N1 free from x)"dy=c"


"\int (2xe^y+2x/y+1/y^3)dx=c"


"x^2e^y+x^2\/y+x\/y^3=c"


4.

"\frac{\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x}}{M}=-tany"

I.f.:

"e^{\int-tanydy}=cosy"


Then:

"(y\/x secy-tany)cosydx+(secy logx-x)cosydy=0"

"(y\/x-sinx)dx-(xcosy-logx)dy=0"

"\int (y/x-sinx)dx=ylogx-xsiny"


Solution:

"ylogx-xsiny=c"


5.

"\frac{3yy'}{y^2-4}=x"


"\int\frac{3ydy}{y^2-4}=\int xdx"


"3ln(y^2-4)=x^2+c_1"

"y=\sqrt{ce^{x^2/3}+4}"